# IOI: Invisible One-Iteration Adversarial Attack on No-Reference Image- and Video-Quality Metrics

### v1

@@ Appendix E Video sequences
Appendix F Subjective experiment setup
To derive subjective scores for adversarial videos, we conducted a crowd-sourced subjective comparison on the Subjectify.us service (sub, Accessed: Jan 2024). For comparison, we compressed all videos, including the original ones, using the x264 video codec with a CRF value of 16 (preset “Medium”). Each pair shown to participants consisted of two samples of the same source video attacked by various attack methods. Each participant was presented with a random pair of videos sequentially and was asked to choose the video with the superior visual quality. An option “Can’t choose” was also provided. Videos were pre-loaded in the browser to prevent delays in playback, and participants had the flexibility to replay the videos multiple times. Each participant compared 12 video pairs, of which two were for verification. Answers from 200 participants who failed the verification were excluded. We collected 8220 responses from 685 successful participants and calculated subjective scores using the Bradley-Terry model (Bradley & Terry, 1952). The average payment to crowdworkers per pair of sequences was \$0.05. We estimate the overall cost of the subjective tests was \$410. Figure 7 presents the subjective experiment’s general process.
@@ Appendix F Subjective experiment setup
Appendix G Per-video results
Tables 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22 contain results of subjective comparison of proposed IOI adversarial attack with eight prior attacks for each video.
Appendix H Attack examples on images and videos
IOI adversarial images and videos are available in the Supplementary materials zip archive. They can also be found at link is hidden for a blind review.

### v2

@@ Appendix I IOI attack on complex VQA metric
Appendix K Subjective experiment setup
To derive subjective scores for adversarial videos, we conducted a crowd-sourced subjective comparison on the Subjectify.us service (sub, Accessed: Jan 2024).
For comparison, we compressed all videos, including the original ones, using the x264 video codec with a CRF value of 16 (preset “Medium”). Each pair shown to participants consisted of two samples of the same source video attacked by various attack methods. Each participant was presented with a random pair of videos sequentially and was asked to choose the video with the superior visual quality. An option “Can’t choose” was also provided. Videos were pre-loaded in the browser to prevent delays in playback, and participants had the flexibility to replay the videos multiple times. Each participant compared 12 video pairs, of which two were for verification. Answers from 200 participants who failed the verification were excluded.
We collected 8220 responses from 685 successful participants and calculated subjective scores using the Bradley-Terry model (Bradley & Terry, 1952). The average payment to crowdworkers per a pair of sequences was \$0.05. We estimate the overall cost of the subjective tests was \$410. Figure 7 presents the subjective experiment’s general process.
@@
Appendix L Per-video results
Tables 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25 contain results of subjective comparison of proposed IOI adversarial attack with eight prior attacks for each video.
Appendix M Attack examples on images and videos
IOI adversarial images and videos are available in the zip archive: https://drive.google.com/file/d/1nrvV70Q4W0vh-2FdWrXHUhMDzYcI6zY1/view?usp=sharing.